King Fisher Aviation has projected the following quarterly sales amounts for the coming year: Quarter 1 $750 Quarter 2 $820 Quarter 3 $790 Quarter 4 $950 Accounts Receivable at the beginning of the year are $350. King Fisher has a 30-day collection period. Calculate the cash collections in each of the four quarters by completing the following for each quarter Beginning receivables Sales Cash Collections Ending Receivables Rework the Ending Receivables calculation using 45 and 60 days What do you recommend to King Fisher regarding changing their 30-day collection period? Explain your reasoning.
Input Area: Beginning A/R a. Collection period b. Collection period S 350 30 45 60 c. Collection period Q3 Q1 Q2 Q4 Sales 750 820 S 790 950 Output Area: 30 -day collection period Q1 Q2 Q3 Q4 Beginning receivables Sales Cash collections Ending receivables 45 Beginning receivables Sales Cash collections Ending receivables 60 C Beginning receivables Sales Cash collections Ending receivables
Answer #1

For quarter one

Beginning account receivable 350

Sales 750

Cash collections = 750× 90/120 + 350 = 912.5

Ending receivables

For 30 days = 750 × 30/120= 187.5

For 45 days = 750 × 45/120 = 281.25

For 60 days = 750 × 60/120 = 375

For quarter 2

Beginning account receivable 187.5

Sales 820

Cash collections = 820× 90/120 + 187.5 = 802.5

Ending receivables

For 30 days = 820 × 30/120= 205

For 45 days = 820× 45/120 = 307.5

For 60 days = 820 × 60/120 = 410

For quarter three

Beginning account receivable 205

Sales 790

Cash collections = 790× 90/120 + 205 = 797.5

Ending receivables

For 30 days = 790 × 30/120= 197.5

For 45 days = 790 × 45/120 = 296.25

For 60 days = 790 × 60/120 = 395

For quarter four

Beginning account receivable 197.5

Sales 950

Cash collections = 950× 90/120 + 197.5 = 900

Ending receivables

For 30 days = 950 × 30/120= 237.5

For 45 days = 950 × 45/120 = 356.25

For 60 days = 950 × 60/120 = 475